OpenStudy (appleduardo):

whats the integral for dx/x^2 + 16 ?

OpenStudy (anonymous):

Trig sub perhaps?

OpenStudy (appleduardo):

i thinks its \[\frac{ \ln (x ^{2} +16)}{ 2 }+C\]

OpenStudy (anonymous):

\[ x=4\tan\theta \]

OpenStudy (anonymous):

This is it right?\[ \int \frac{dx}{x^2+16} \]

OpenStudy (appleduardo):

yep!

OpenStudy (anonymous):

Wanna do a trig sub. You can't do the \(\ln(\;)\) thing unless the denominator is linear.

OpenStudy (anonymous):

\[ x=4\tan\theta,\,dx=4\sec^2\theta \,d\theta \]

OpenStudy (anonymous):

Try this substitution and see what happens.

OpenStudy (appleduardo):

i'll doit, thanks!